OpenStudy (anonymous):

what is the solution set for the equation x^2+2x=15 answer please

OpenStudy (anonymous):

x^2+2x=15 x^2+2x-15=0 (x+5)(x-3)=0 x+5=0 x-3=0 x=-5 x=3

OpenStudy (anonymous):

thanks